find the distance A(2,3) measured parallel to the line x-y=5 from the line 2x+y+6=0
 

Grade:11

1 Answers

Muna Kunwar
7 years ago
given equation x-y=5 and 2x+y+6=0 x=5+y..... eq1 2(5+y)+y+6 10+2y+y+6=0 y=-16/3 and x=-1/3 Now, distance between A and (x,y)is root under (x2-x1) +(y2-y1)2 hence distance =8.05